OpenStudy (anonymous):

Complete the ordered pairs so that each solution is given linear equation. x=6-y x is 0, 3, 6 and y is ?,?,? Than graph equation

OpenStudy (anonymous):

Plug in each of the numbers into the equation to get the y value. (0,_) (3,_)(6,_)

OpenStudy (anonymous):

ok so x=6-0=6 x=6-3= -3 x=6-6=-0

OpenStudy (anonymous):

(0,6), (3,-3), (6,-0)

OpenStudy (anonymous):

Check your math on x-value 3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

3 = 6-y. Get y on its own. y = 6-3

OpenStudy (anonymous):

y=6-3 =3

OpenStudy (anonymous):

is my ordered pair wrong

OpenStudy (anonymous):

its (3,3) not (3,-3)

OpenStudy (anonymous):

so (0,6), (3,3),(6,6)

OpenStudy (anonymous):

Looks right to me.

OpenStudy (anonymous):

ok so i have to plot graph

OpenStudy (anonymous):

how do i plot graph because i have two parralel lines

OpenStudy (anonymous):

you have two parallel lines where?

OpenStudy (anonymous):

on my graph

OpenStudy (anonymous):

Like. two random lines?

OpenStudy (anonymous):

well i plot 0,6 as a scatter line and the other two paris lines wher two parallel lines like this

OpenStudy (anonymous):